# Functional Stochastic Gradient MCMC for Bayesian Neural Networks

### v1

@@ D.1 The effects of the sample size for naive SGMCMC and functional SGMCMC
D.2 The effects of the number of measurement points on the gradient estimation of functional prior
@@
D.3 The effects of the kernel function on the functional GP prior
(a) fSGLD-RBF
(b) fSGLD-Linear
(c) fSGLD-Matern
(d) fSGHMC-RBF
(e) fSGHMC-Linear
(f) fSGHMC-Matern
@@ D.3 The effects of the kernel function on the functional GP prior
Functional Gaussian processes (GPs) priors are able to encode prior knowledge about function properties (e.g., periodicity and smoothness) through corresponding kernel functions. We used RBF kernel in the 1-D extrapolation experiment, which is suited for modeling polynomial functions. We now consider two other kernel functions of GP prior to our functional SGMCMC: Matern kernel and Linear kernel (unsuitable for modeling polynomial oscillatory curves). The results are shown in Figure 5. The results from the Linear kernel in Figure 5(b) and Figure 5(e) for fSGLD and fSGHMC respectively show that the fitting performance deteriorates significantly in the observation region and also demonstrate a strong linear trend (converging to a horizontal line) in the middle unseen region $[-0.25,0.25]$. The uncertainty intervals in the leftmost $[-1,-0.75]$ and rightmost $[0.75,1]$ non-observation regions deviate far from the true function. On the contrary, the results of the Matern kernel are hardly different from those of the RBF kernel, due to the fact that it also has a strong ability to describe polynomial functions.

### v2

@@ H.1 The effects of the sample size for naive SGMCMC and functional SGMCMC
H.2 The effects of the number of measurement points on the gradient estimation of functional prior
@@
H.3 The effects of the functional GP prior
(a) GP prior-RBF
(b) GP prior-Linear
(c) GP prior-Matern
(d) fSGLD-RBF
(e) fSGLD-Linear
(f) fSGLD-Matern
(g) fSGHMC-RBF
(h) fSGHMC-Linear
(i) fSGHMC-Matern
@@ H.3 The effects of the functional GP prior
Functional Gaussian processes (GPs) priors are able to encode prior knowledge about function properties (e.g., periodicity and smoothness) through corresponding kernel functions. We used RBF kernel in the 1-D extrapolation experiment, which is suited for modelling polynomial functions. We now consider two other kernel functions of GP priors to our functional SGMCMC: Matern kernel and Linear kernel (unsuitable for modelling polynomial oscillatory curves). The results are shown in Figure 6. The top row presents the GP prior predictions for RBF kernel, Linear kernel and Matern kernel, respectively. The results from the mismatched Linear kernel in Figure 6(e) and Figure 6(h) for fSGLD and fSGHMC, respectively, show that the fitting performance deteriorates in the observation region and also demonstrate a strong linear trend (converging to a horizontal line) in the middle unseen region $[-0.25,0.25]$ since the Linear GP prior prediction shows a completely linear trend and falls to fit the target function. The uncertainty intervals in the leftmost $[-1,-0.75]$ and rightmost $[0.75,1]$ non-observation regions deviate far from the true function. On the contrary, the results of the Matern kernel are hardly different from those of the RBF kernel due to the fact that it also has a strong ability to describe polynomial functions. This indicates that our method can effectively incorporate functional prior information into the posterior inference.
We also plot the predictions of fSGLD and fSGHMC with no pre-trained GP priors in Figure 7. The leftmost columns, Figure 7(a) and Figure 7(c), show the results for fSGLD and fSGHMC with pre-trained GP priors, as described in the main paper. The rightmost columns, Figure 7(b) and Figure 7(d), display the predictions without pre-trained GP priors, using a fixed constant mean function $m(x)=0$ and a kernel function $k(x,x^{\prime})=1$. Despite not having pre-trained priors, the posterior results remain quite similar to those with pre-trained GP priors, indicating that our methods are robust to the absence of prior pre-training. In Table 12, we report the RMSE and NLL results of fSGLD and fSGHMC on the UCI regression (Yacht) dataset without pre-trained GP priors. Even though the fixed GP priors without pre-training show slight degradation, the performance remains competitive compared to other baseline methods.
(a) fSGLD
(b) fSGLD
(c) fSGHMC
(d) fSGHMC
Figure 7: 1-D extrapolation example without prior pre-training. (a) and (c) are the results with pre-trained prior; (b) and (d) are results without pre-trained prior.
Table 12: RMSE and NLL of fSGLD and fSGHMC on UCI regression (yacht) with no pre-trained GP priors.
Model pre-trained GP fixed GP
RMSE
fSGLD 0.41 0.58
fSGHMC 0.25 0.51
NLL
fSGLD -2.46 -1.65
fSGHMC -3.37 -2.52
Furthermore, we analyze the influence of different pre-training epochs for the GP priors by comparing results with 100, 20, and 10 pre-training epochs. Figure 8 provides a detailed comparison: the leftmost columns, Figure 8(a), Figure 8(d) and Figure 8(g) depict the predictions of the three different GP priors, while the middle and rightmost columns illustrate the fSGLD and fSGHMC posterior results from these priors, respectively. Interestingly, while the predictions from the three GP priors differ significantly, there is no noticeable difference in the corresponding fSGLD and fSGHMC posteriors. This observation highlights that the effectiveness of our methods does not strongly depend on the GP prior pre-training procedure, making them more flexible in practical applications.
(a) GP-100 epochs
(b) fSGLD
(c) fSGHMC
(d) GP-20 epochs
(e) fSGLD
(f) fSGHMC
(g) GP-10 epochs
(h) fSGLD
(i) fSGHMC
Figure 8: 1-D extrapolation example of different pre-training epochs for GP prior. The leftmost column shows the predictions of GP prior with three different pre-training epochs: 100, 20, and 10 epochs from top to bottom. The middle and the rightmost columns are the corresponding fSGLD and fSGHMC posteriors from these three GP prior, respectively.